5. In a construction site, two vertical cables are supporting a 3.00m steel beam weighing 1500N. Suppose cable A can support a maximum weight of 2000N while cable B can support a maximum weight of 3000N. A construction worker would like to place an object on the horizontal beam, what is the maximum mass that he can put without breaking the cables? where should the object be located?
1
Expert's answer
2021-10-13T16:56:18-0400


Write equilibrium of forces:


"R_2+R_1-mg-Mg=0,\\
2000+3000-1500-Mg=0,\\
Mg=3500\text{ N}."

Equilibrium of torques (around point where R2 starts):


"0=-1.5·1500-(1.5+x)·3500+3000·3,\\
x=0.429\text{ m}."